# Hyperparameter Transfer with Mixture-of-Experts Layers

### v1

@@ 1 Introduction
This motivates the study of HP transfer (Yang & Hu, 2022; Yang et al., 2022; Bordelon & Pehlevan, 2022; Bordelon et al., 2023; Yang et al., 2023b; Bordelon & Pehlevan, 2025; Bordelon et al., 2024; Dey et al., 2025; Bergsma et al., 2025a; Wang & Aitchison, 2025), a suite of techniques for obtaining performant HPs in large models directly from good HPs found by tuning substantially smaller models. This HP extrapolation, or transfer, is determined through a parameterization, a set of theoretically-motivated rules that predict how variations in each scaling dimension of interest (e.g. model width, depth) should modify raw HP values.
@@
Theoretical grounding.  We provide a theoretical justification of our proposed parameterization using dynamical mean-field theory (DMFT) (Bordelon & Pehlevan, 2022). Specifically, we obtain an explicit description of the training dynamics of residual networks with MoE layers in the simultaneous limit of infinite width, depth, expert size, and expert count. The evolution of (finite) network summary statistics (e.g. layerwise feature kernels) is therefore automatically consistent across scale. We find a novel three-level mean-field hierarchy: residual stream representations are mean-field over expert outputs, which are themselves mean-field over individual expert neurons. The DMFT analysis justifies several subtle but important properties of our parameterization, such as the transfer of optimal HPs across expert hidden width multiplier (Figure 2, last column).
@@
Insights on architecture shapes.  We empirically verify, via our zero-shot optimal HP, existing findings (Krajewski et al., 2024; Boix-Adsera & Rigollet, 2025) on scaling expert size versus expert count in MoEs. Without the need to retune HPs at each scale, we consistently recover the benefit of increasing number of experts over expert sizes at fixed parameter count (Figure˜5 and Figure˜16).
Figure 1: Matching active architecture to be GPT2-small (124M) and 500K batch size, comparison of MoE training loss on our zero-shot Adam HPs (found from tuning 38M activated base models) on FineWeb versus (dense GPT) baseline (Karpathy, 2023), speedrun AdamW, and speedrun Muon loss curves (Jordan & contributors, 2025) towards the 3.28 val. loss for baseline at 10B tokens. See Section˜5 for details. Dense benchmarks are taken before Oct.14, 24, with advanced mods such as zero down-proj. and QK-norms.
@@ 2 Related backgrounds
Figure 2: Global base learning rate (first row) and global base init (second row) transfer trained on 1B tokens (2000 steps) on the Fineweb dataset, with different model sizes from 20M to 1.8B scaling across width, depth, number of experts (fixing sparsity), and expert MLP hidden multiplier. We fix the base config ($n_{\operatorname{embd}}\text{ (W) }=512$, $L=8$, $(n_{\operatorname{exp}},n_{\operatorname{act}})=(4,1)$ (‘E4A1’ in the figure), $\alpha=1$) and vary one dimension at a time. Error bars in the last row are (max, min, and median) over four independent seeds. See Section 5 for details.
@@
HP transfer and DMFT.  Traditional approaches to HP tuning at large scale require either grid-searching in a large model or extrapolating from power laws fit to a family of smaller models (Li et al., 2025; Zhou et al., 2026). Instead, our approach seeks the direct (rule-based) transfer of optimal HP from small models to larger ones. The core behind transfer is parameterizations that stabilize the forward and backward passes across scales. First studied for scaling the width of MLPs, this stability condition resulted in two types of parameterizations: the Neural Tangent parameterization (NTP) (Hayou et al., 2022; Jacot et al., 2020; Roberts et al., 2022) and the Mean-Field parameterization (MFP) (Mei et al., 2018; Chizat et al., 2019; Rotskoff & Vanden-Eijnden, 2022). Such stability conditions have also been studied via the modular norm (Large et al., 2024) and effective field theory (Dinan et al., 2023; Yaida, 2022; Roberts et al., 2022).
@@
Dynamical mean-field theory (DMFT) provides a way to theoretically analyze HP transfer by showing which parameterizations cause training dynamics of finite networks to converge to a well-defined limit. This framework traces the evolution (in infinite-sized networks) of mean-field model statistics, such as inner-product kernels for hidden activations and gradients. DMFT results were first derived for MLPs (Bordelon & Pehlevan, 2022) and later extended for multi-head self-attention (Bordelon et al., 2024), transformers in the infinite-depth limit (Bordelon et al., 2023), and finite-width correction (Bordelon & Pehlevan, 2025). We refer to Sec.4 in (Bordelon et al., 2023) for a more thorough explanation of how DMFT backs up HP transfer.
@@ 3.1 Model setup
Our theory and experiments concern decoder-only Transformer language models (Radford et al., 2019) with MoE modules in the feed-forward (FFN) layers. In this architecture, a sequence of $T$ input tokens is mapped to an output by first up-projecting (along with positional embedding) each token to a sequence $h^{(0)}$ in $\mathbb{R}^{n_{\operatorname{embd}}\times T}$. This initializes the residual stream, whose updates are computed through $2L$ residual layers that intersperse $L$ pre-LayerNorm multi-head self-attention (MHSA) blocks (for $\ell=1,\ldots,L$):
@@
where $g_{i}(h)\triangleq\sigma(r_{i})\in\mathbb{R}$ are (un-normalized) mixing coefficients, $\sigma$ is a non-linear activation function, and
@@
with a hidden layer of size $\alpha_{\operatorname{ffn}}\cdot n_{\operatorname{embd}}\in\Omega(n_{\operatorname{embd}})$, and $W^{(i)}_{\operatorname{up}},W^{(i)}_{\operatorname{down}}\in\mathbb{R}^{n_{\operatorname{embd}}\times\alpha_{\operatorname{ffn}}n_{\operatorname{embd}}}$. To be close to our theoretical analysis (Section˜4), our setup disentangles inter-expert router weights (following (Wang et al., 2024)): for each expert $i$, router $g_{i}(h)$ only depends on $W_{\operatorname{router}}^{(i)}\in\mathbb{R}^{n_{\operatorname{embd}}}$. Practically, we did not observe significant differences in model performance across scale between different types of mixing coefficients (e.g., topK-then-softmax, softmax-then-topK).
Training. We consider the standard Adam optimizer (Kingma & Ba, 2017) (see Apdx. C for discussion on weight decay). During training, we treat the activated experts sets $A(h)$ in each layer as no-grad vectors, and router matrices only receive gradients through the expert mixing coefficients $g_{i}(h)$. Writing $\operatorname{Load}_{i}\in[0,1]$ for the (batch) proportion of tokens routed to expert $E_{i}$, we encourage expert load-balancing, i.e. we ask that
@@ 3.2 Scaling that preserves sparsity versus topK
We have three practical motivations for this choice. First, (assuming perfect balancing) in a batch of $B$ tokens, each MoE expert sees only $\kappa B$ tokens, whereas attention and routing modules see the full $B$ tokens. This suggests HPs will not transfer across different $\kappa$ (this is partly confirmed in Figure 11). Second, for large deployments, it has been empirically reported that the range of optimal sparsity $\kappa$ is often hardware-controlled (StepFun et al., 2025) and bounded below. Finally, a constant $\kappa$ is natural in the mean-field analysis (Sec. 4) as it represents conditioning on a constant-probability event when integrating over the mean-field measure of experts. Our theory suggests that HPs will transfer at fixed $\kappa$ and our empirical results support this point.
@@ 3.3 MoE Parametrization
Notation. For any vector $\theta\in\mathbb{R}^{k}$ we will shorthand $\theta\in\Theta(C)$ to denote $\|\theta\|_{2}^{2}\in\Theta(k\cdot C)$. To derive scaling rules for Adam, we will also use $\overline{\nabla W}$ to denote the normalized Adam update per step $W_{t+1}\leftarrow W_{t}-\eta\overline{\nabla W_{t}}$ and use Adam’s approximation to SignGD ($\overline{\nabla W}\in\{-1,1\}^{*}$) (Bernstein et al., 2018) so $\Delta w=\eta\overline{\nabla w}\approx\eta\cdot\operatorname{sgn}\left(\frac{\partial w}{\partial\Theta}\right)$ for any weight group $w$. We also use $\cos(v_{1},v_{2})=v_{1}^{T}v_{2}/\|v_{1}\|_{2}\|v_{2}\|_{2}$ to denote the cosine of angle between two vectors.
@@
Router weights. We start with deriving the parameterization for the MoE router matrix $W_{\operatorname{router}}$. We apply (3) on $W=W_{\operatorname{router}}^{(i)}$ and $z=g_{i}(h)=\sigma\left(h^{T}\cdot W_{\operatorname{router}}^{(i)}\right)$. Assuming that $\sigma^{\prime}(\cdot)\in\Theta(1)$, we arrive at our first desiderata:
@@
Recall $\Delta W_{\operatorname{router}}^{(i)}=-\eta_{\operatorname{router}}\overline{\nabla W_{\operatorname{router}}^{(i)}}.$ We assume
$\frac{{h^{T}\cdot\overline{\nabla W_{\operatorname{router}}^{(i)}}}{}}{\|\overline{\nabla W_{\operatorname{router}}}\|_{2}\|h\|_{2}}=\cos\left(\,\overline{\nabla W_{\operatorname{router}}^{(i)}}\,,\,h\right)\in\Theta(1),$
@@
We defer the derivation of $\eta_{\operatorname{bias}}$ to Apdx. B. As before, max-update principles place no requirement on initialization so long as $b\in\Theta(1)$. However, taking into account expert load balancing, we conveniently initialize biases at zero so no one expert receives dominant tokens (or no tokens) at init.
@@
[…]
@@
Assume for all individual neurons $j\in[\alpha_{\operatorname{ffn}}n_{\operatorname{embd}}]$:
$\frac{h^{T}(\Delta W_{\operatorname{up}})_{j}}{\|\Delta(W_{\operatorname{up}})_{j}\|_{2}\|h\|_{2}}=\cos\left(\Delta(W_{\operatorname{up}})^{T}_{j},h\right)=\Theta(1)$
which dictates LR for the up-projection via requiring that $\|\Delta(W_{\operatorname{up}})_{j}\|_{2}\in\Theta(\|h\|_{2}^{-1})=\Theta(n_{\operatorname{embd}}^{-1/2})$. For $\sigma_{\operatorname{init}}$,
$\Theta(\sqrt{\alpha_{\operatorname{ffn}}n_{\operatorname{embd}}})\ni\|(W_{\operatorname{up}})^{T}\Delta h\|_{2}$ $\approx\|\Delta h\|_{2}\|W_{\operatorname{up}}\|_{\operatorname{op}}$
and given that $\|\Delta h\|_{2}\approx\sqrt{n_{\operatorname{embd}}}$ we get $\|W_{\operatorname{up}}\|_{\operatorname{op}}\in\Theta(\sqrt{\alpha_{\operatorname{ffn}}})$. We can thus conclude following standard results from random matrix theory $\|W_{\operatorname{up}}\|_{\operatorname{op}}\approx\sqrt{\alpha_{\operatorname{ffn}}n_{\operatorname{embd}}}\sigma_{\operatorname{init}}$ when assuming $\alpha_{\operatorname{ffn}}\in\Omega(1)$.
For the down-projection layers, similar derivations can be made for the learning rate, where we assume that the up-projected embedding is correlated with weight updates $\operatorname{cos}(\Delta(W_{\operatorname{down}})_{\cdot},\phi(h_{\operatorname{up}}))=\Theta(1)$ across all neurons. For the initialization scale, our conditions require the following norm to be in $\Theta(n_{\operatorname{embd}}^{1/2})$:
@@
$\approx\sqrt{\alpha_{\operatorname{ffn}}n_{\operatorname{embd}}}\sqrt{\alpha_{\operatorname{ffn}}n_{\operatorname{embd}}}\sigma_{\operatorname{init}}(W_{\operatorname{down}})$
when $\alpha_{\operatorname{ffn}}\in\Omega(1)$. Combined with the above, we have:
@@ 4 Dynamical mean field theory
To theoretically establish that our parameterization enables a scale-invariant well-defined limit (and consequently admits HP transfer across jointly scaling model dimensions), we study the training dynamics for a deep residual MoE model with diverging residual stream width and expert count $n_{\operatorname{embd}},n_{\operatorname{exp}}\to\infty$ with activated experts $n_{\operatorname{act}}=\kappa n_{\operatorname{exp}}$. We can also optionally take the expert size $n_{\operatorname{hid}}\triangleq\alpha_{\operatorname{ffn}}n_{\operatorname{embd}}\to\infty$ and the depth $L\to\infty$ so long as ${n_{\operatorname{embd}}}/({n_{\operatorname{exp}}n_{\operatorname{hid}}L)}$ is bounded in the limit. We find that the MoE dynamics have a three-level mean-field structure [^1] By saying “$X$ is mean-field”, we mean the dynamics of each individual element of the population depend only on itself and averages over the entire population. See Apdx E for details.:
@@
    The network outputs are determined by population averages over residual stream neurons and averages over the state of the experts within each hidden MoE layer. Residual stream neurons are mean-field in the $n_{\operatorname{embd}},n_{\operatorname{exp}}\to\infty$ limit. The macroscopic behavior of the network is summarized by averages (denoted by $\left$$\cdot\right$>) over the residual stream neuron distribution.
@@
    Each expert is characterized by a collection of scalar variables that include the router pre-activation and bias. These expert state variables are mean-field in the $n_{\operatorname{embd}},n_{\operatorname{exp}}\to\infty$ limit, with averaging denoted as $\left[\cdot\right]$. The top-$n_{\operatorname{act}}$ action is determined by expert random variables $q=\sigma(r)+b$ and hard gating is set by a quantile threshold $q_{\star}(\kappa)$ which satisfies $\left[\mathbf{1}_{q\geq q_{\star}(\kappa)}\right]=\kappa$. The (hard) gating variables for each expert are thus $\sigma\triangleq\mathbf{1}_{q\geq q_{\star}(\kappa)}\sigma(r)$.
@@
    Within each expert, hidden neuron states are mean-field in the $n_{\operatorname{hid}}\to\infty$ limit. Their dynamics depend on all three levels of the mean-field hierarchy. We denote averaging over the within-expert neurons $\left\{\cdot\right\}$.
At a high level, the training dynamics close in terms of a finite set of deterministic quantities $\bm{Q}$ which include correlation and response kernels, as well as the network output. Some examples of the relevant correlation kernels include
$C^{\ell}_{h}(x,x^{\prime},t,t^{\prime})=\left<h^{\ell}(x,t)h^{\ell}(x^{\prime},t^{\prime})\right$>
$M^{\ell}_{\sigma\sigma C_{\phi_{k}}}(x,x^{\prime},t,t^{\prime})$
$=\left[\sigma^{\ell}(x,t)\sigma^{\ell}(x^{\prime},t^{\prime})\{\phi(h^{\ell}(x,t))\phi(h^{\ell}(x^{\prime},t^{\prime}))\}\right].$ (4)
for input $x$ at time $t$. We derive the full set of closed evolution equations for the macroscopic variables $\bm{Q}$ that govern the network outputs in Appendix E. Some useful observations can be immediately gleaned from the theory:
@@
    $n_{\operatorname{hid}}$ does not need to diverge for network outputs and residual stream measures to concentrate. If $n_{\operatorname{embd}},n_{\operatorname{exp}}\to\infty$ with $n_{\operatorname{exp}}/n_{\operatorname{embd}}=\nu$, the output of the network obeys a deterministic evolution which depends on $n_{\operatorname{hid}}$ explicitly, but not on $\nu$, similar to dependence on head size in mean-field limits of attention blocks (Hron et al., 2020; Bordelon et al., 2024). This MoE regime is empirically studied in (He, 2024).
@@
A similar three-level structure also appears in multi-head self-attention, where there is a measure over attention heads (analogous to the expert measure) and an additional measure of key-query neurons within each head (Bordelon et al., 2024). Combining our novel analysis of the MoE layers with the existing DMFT for multi-head attention yields a complete description of sparse transformers in this work. Further details are deferred to Appendix E.

### v2

@@ 1 Introduction
This motivates the study of HP transfer (Yang and Hu, 2022; Yang et al., 2022; Bordelon and Pehlevan, 2022; Bordelon et al., 2023; Yang et al., 2023b; Bordelon and Pehlevan, 2025; Bordelon et al., 2024; Dey et al., 2025; Bergsma et al., 2025a; Wang and Aitchison, 2025), a suite of techniques for obtaining performant HPs in large models directly from good HPs found by tuning substantially smaller models. This HP extrapolation, or transfer, is determined through a parameterization, a set of theoretically-motivated rules that predict how variations in each scaling dimension of interest (e.g. model width, depth) should modify raw HP values.
@@
Theoretical grounding.  We provide a theoretical justification of our proposed parameterization using dynamical mean-field theory (DMFT) (Bordelon and Pehlevan, 2022). Specifically, we obtain an explicit description of the training dynamics of residual networks with MoE layers in the simultaneous limit of infinite width, depth, expert size, and expert count. The evolution of (finite) network summary statistics (e.g. layerwise feature kernels) is therefore automatically consistent across scale. We find a novel three-level mean-field hierarchy: residual stream representations are mean-field over expert outputs, which are themselves mean-field over individual expert neurons. The DMFT analysis justifies several subtle but important properties of our parameterization, such as the transfer of optimal HPs across expert hidden width multiplier (Figure 2, last column).
@@
Insights on architecture shapes.  We empirically verify, via our zero-shot optimal HP, existing findings (Krajewski et al., 2024; Boix-Adsera and Rigollet, 2025) on scaling expert size versus expert count in MoEs. Without the need to retune HPs at each scale, we consistently recover the benefit of increasing number of experts over expert sizes at fixed parameter count (Figure˜5 and Figure˜16).
Figure 1: Matching active architecture to be GPT2-small (124M) and 500K batch size, comparison of MoE training loss on our zero-shot Adam HPs (found from tuning 38M activated base models) on FineWeb versus (dense GPT) baseline (Karpathy, 2023), speedrun AdamW, and speedrun Muon loss curves (Jordan and contributors, 2025) towards the 3.28 val. loss for baseline at 10B tokens. See Section˜5 for details. Dense benchmarks are taken before Oct.14, 24, where more advanced architecture modifications such as zero down-projection init. and QK-norms are applied.
@@ 2 Related backgrounds
Figure 2: Global base learning rate (first row) and global base init (second row) transfer trained on 1B tokens (2000 steps) on the Fineweb dataset, with different model sizes from 20M to 1.8B scaling across width, depth, number of experts (fixing sparsity), and expert MLP hidden multiplier. We fix the base config ($n_{\operatorname{embd}}\text{ (W) }=512$, $L=8$, $(n_{\operatorname{exp}},n_{\operatorname{act}})=(4,1)$ (‘E4A1’ in the figure), $\alpha_{\operatorname{ffn}}=1$) and vary one dimension at a time. Error bars in the last row are (max, min, median) over four independent seeds. See Sec. 5 for details.
@@
HP transfer and DMFT.  Traditional approaches to HP tuning at large scale require either grid-searching in a large model or extrapolating from power laws fit to a family of smaller models (Li et al., 2025; Zhou et al., 2026). Instead, our approach seeks the direct (rule-based) transfer of optimal HP from small models to larger ones. The core behind transfer is parameterizations that stabilize the forward and backward passes across scales. First studied for scaling the width of MLPs, this stability condition resulted in two types of parameterizations: the Neural Tangent parameterization (NTP) (Hayou et al., 2022; Jacot et al., 2020; Roberts et al., 2022) and the Mean-Field parameterization (MFP) (Mei et al., 2018; Chizat et al., 2019; Rotskoff and Vanden-Eijnden, 2022). Such stability conditions have also been studied via the modular norm (Large et al., 2024) and effective field theory (Dinan et al., 2023; Yaida, 2022; Roberts et al., 2022).
@@
Dynamical mean-field theory (DMFT) provides a way to theoretically analyze HP transfer by showing which parameterizations cause training dynamics of finite networks to converge to a well-defined limit. This framework traces the evolution (in infinite-sized networks) of mean-field model statistics, such as inner-product kernels for hidden activations and gradients. DMFT results were first derived for MLPs (Bordelon and Pehlevan, 2022) and later extended for multi-head self-attention (Bordelon et al., 2024), transformers in the infinite-depth limit (Bordelon et al., 2023), and finite-width correction (Bordelon and Pehlevan, 2025).
@@ 3.1 Model setup
Our theory and experiments concern decoder-only Transformer language models (Radford et al., 2019) with MoE modules in the feed-forward (FFN) layers. In this architecture, a sequence of $T$ input tokens is mapped to an output by first up-projecting (along with positional embedding) each token to a sequence $h^{(0)}$ in $\mathbb{R}^{n_{\operatorname{embd}}\times T}$. This initializes the residual stream, whose updates are computed through $2L$ residual layers that intersperse $L$ pre-LayerNorm multi-head self-attention (MHSA) blocks (for $\ell=0,\ldots,L-1$):
@@
where $g_{i}(h)\triangleq\sigma(r_{i})\in\mathbb{R}$ are (un-normalized) mixing coefficients, $\sigma$ is a non-linear (sigmoid) function, and
@@
with a hidden layer of size $\alpha_{\operatorname{ffn}}\cdot n_{\operatorname{embd}}\in\Omega(n_{\operatorname{embd}})$, and $W^{(i)}_{\operatorname{up}},W^{(i)}_{\operatorname{down}}\in\mathbb{R}^{n_{\operatorname{embd}}\times\alpha_{\operatorname{ffn}}n_{\operatorname{embd}}}$. To be close to our theoretical analysis (Section˜4), our setup disentangles mixing coefficients for different experts (following (Wang et al., 2024; ArceeAI et al., 2026)): for each expert $i$, router $g_{i}(h)$ only depends on $W_{\operatorname{router}}^{(i)}\in\mathbb{R}^{n_{\operatorname{embd}}}$. Practically, we did not observe significant differences in model performance across scale between different types of mixing coefficients (e.g., topK-then-softmax, softmax-then-topK).
Training. We consider the standard Adam optimizer (Kingma and Ba, 2017) (see Apdx. C for discussion on weight decay). During training, we treat the activated experts sets $A(h)$ in each layer as no-grad vectors, and router matrices only receive gradients through the expert mixing coefficients $g_{i}(h)$. Writing $\operatorname{Load}_{i}\in[0,1]$ for the (batch) proportion of tokens routed to expert $E_{i}$, we encourage expert load-balancing, i.e. we ask that
@@ 3.2 Scaling that preserves sparsity versus topK
We have three practical motivations for this choice. First, (assuming perfect balancing) in a batch of $B$ tokens, each MoE expert sees only $\kappa B$ tokens, whereas attention and routing modules see the full $B$ tokens. This suggests HPs will not transfer across different $\kappa$ (this is partly confirmed in Figure 11). Second, for large deployments, it has been empirically reported that the range of optimal sparsity $\kappa$ is often hardware-controlled (StepFun and others, 2025) and bounded below. Finally, a constant $\kappa$ is natural in the mean-field analysis (Sec. 4) as it represents conditioning on a constant-probability event when integrating over the mean-field measure of experts. Our theory suggests that HPs will transfer at fixed $\kappa$ and our empirical results support this point.
@@ 3.3 Proposed MoE Parametrization
Notation. For any vector $\theta\in\mathbb{R}^{k}$ we will shorthand $\theta\in\Theta(C)$ to denote $\|\theta\|_{2}^{2}\in\Theta(k\cdot C)$. To derive scaling rules for Adam, we will also use $\overline{\nabla W}$ to denote the normalized Adam update per step $W_{t+1}\leftarrow W_{t}-\eta\overline{\nabla W_{t}}$ and use Adam’s approximation by SignGD ($\overline{\nabla W}\in\{-1,1\}^{*}$) (Bernstein et al., 2018) so $\Delta w=\eta\overline{\nabla w}\approx\eta\cdot\operatorname{sgn}\left(\frac{\partial\mathcal{L}}{\partial w}\right)$ for any weight group $w$. We also denote by $\cos(v_{1},v_{2})$ the cosine of the angle between two vectors $v_{1},v_{2}$.
@@
To derive from conditions of the form (3) to our parameterization, we will often make so-called “Law of Large Numbers” type alignment assumptions (Yang and Hu, 2022; Yang et al., 2023b; Everett et al., 2024): if $v$ is either a vector of pre-activations or their change after one step of training, and $w$ is either a the vector of model weight updates after one step of training or model weights (respectively), then their dot products scales like:
$v^{T}w/\|v\|\|w\|=\cos(v,w)\in\Theta(1).$ (4)
whereas two random $d$-dimensional vectors typically give alignments of $\cos(v,w)\in O(1/\sqrt{d})$.
Router weights. We start by deriving the parameterization for the MoE router matrix $W_{\operatorname{router}}$. We apply (3) on $W=W_{\operatorname{router}}^{(i)}$ and $z=g_{i}(h)=\sigma\left(h^{T}W_{\operatorname{router}}^{(i)}\right)$. Assuming that $\sigma^{\prime}(\cdot)\in\Theta(1)$, we arrive at our first desiderata:
@@
We defer the heuristic justification of $\eta_{\operatorname{bias}}$ to Appendix B. As before, max-update principles place no requirement on initialization so long as $b\in O(1)$. However, taking into account expert load balancing, we conveniently initialize biases at zero so no one expert disproportionately receives tokens at initialization
@@
Expert MLP weights.  Each individual expert module admits a separate 1-hidden layer MLP. Dropping expert indices, define the notations in a forward pass as
@@
For each individual neuron $j\in\{1,2,\dots,\alpha_{\operatorname{ffn}}n_{\operatorname{embd}}\}$, assuming LLN alignment (4) for $h$ and the row vector $(\Delta W_{\text{up}})_{j}$ sets the up-projection LR in that:
$\eta\|\overline{\nabla(W_{\operatorname{up}})_{j}}\|=\|\Delta(W_{\operatorname{up}})_{j}\|_{2}\in\Theta(\|h\|_{2}^{-1})=\Theta(n_{\operatorname{embd}}^{-1/2}).$
To set $\sigma_{\operatorname{init}}(W_{\operatorname{up}})$ we use a similar alignment assumption
$\|(W_{\operatorname{up}})^{T}\Delta h\|_{2}$ $\approx\|\Delta h\|_{2}\|W_{\operatorname{up}}\|_{\operatorname{op}}\in\Theta(\sqrt{\alpha_{\operatorname{ffn}}n_{\operatorname{embd}}})$
and given that $\|\Delta h\|_{2}\approx\sqrt{n_{\operatorname{embd}}}$, our condition requires $\|W_{\operatorname{up}}\|_{\operatorname{op}}\in\Theta(\sqrt{\alpha_{\operatorname{ffn}}})$. The usual scaling from random matrix theory gives:
$\|W_{\operatorname{up}}\|_{\operatorname{op}}\approx\sqrt{\alpha_{\operatorname{ffn}}n_{\operatorname{embd}}}\sigma_{\operatorname{init}}$
which determines $\sigma_{\mathrm{init}}$. These conditions also guarantees $\Delta h_{\operatorname{up}}\in\Theta(1)$. For $W_{\operatorname{down}}$, the learning rate $\eta(W_{\operatorname{down}})$ is derived similarly via applying (4) to $\left((\Delta W_{\operatorname{down}})_{j},\phi(h_{\operatorname{up}})\right)$, and the initialization is given by
@@
$\approx\sqrt{\alpha_{\operatorname{ffn}}n_{\operatorname{embd}}}\sqrt{\alpha_{\operatorname{ffn}}n_{\operatorname{embd}}}\sigma_{\operatorname{init}}(W_{\operatorname{down}}).$
Asking that to be in $\Theta(n_{\operatorname{embd}}^{1/2})$, we therefore find:
@@
Note that the dependence of $\sigma_{\operatorname{init}}(W_{\operatorname{down}}^{(i)})$ concerning $\alpha_{\operatorname{ffn}}$ is distinct from the standard fan-in initialization (see Figure˜11 for a simple ablation). A heuristic justification (see also (Chizat, 2025)) is to imagine $n_{\operatorname{embd}}=1$ then treat $\alpha_{\operatorname{ffn}}$ as the intermediate “width” in a two-layer MLP under the mean-field parametrization (as opposed to NTP, which yields the standard fan-in initialization).
@@ 4 Dynamical mean field theory
The existence of a well-defined mean-field limit in the training dynamics is a strong justification of HP transfer (Yang and Hu, 2022; Bordelon et al., 2023, 2024). To theoretically establish that our parameterization enables a scale-invariant limit (and consequently admits HP transfer across jointly scaling model dimensions), we outline here a novel approach to studying the training dynamics for a deep residual MoE model [^1] While we focus on analyzing a MoE-block only residual stream for the context of this work, we see no obstruction to deriving complete DMFT equations for the full sparse transformer architecture by combining with existing analysis of multi-head self-attention layers in (Bordelon et al., 2024). in which each input token $x\in\mathbb{R}^{D}$ is processed by first up-projection $h^{(0)}=W_{\mathrm{embed}}x\in\mathbb{R}^{n_{\mathrm{embd}}}$, then passing through $L$ residual MoE feedforward layers
$h^{(\ell+1)}=h^{(\ell)}+L^{-1}f_{\mathrm{MoE}}^{\ell}(h^{(\ell)})\in\mathbb{R}^{n_{\mathrm{embd}}},\quad\ell=1,\ldots,L$
and finally outputting a scalar via the final un-embedding layer $f(x)=W_{\operatorname{unembd}}\phi(h^{(L)})\in\mathbb{R}.$
Using the analog of our parameterization derived in Section˜3.3 for this reduced model, we obtain in Appendix E an explicit mean-field description for the full training dynamics of the model under gradient flow in the limit of diverging residual stream width and expert count $n_{\operatorname{embd}},n_{\operatorname{exp}}\to\infty$ and constant proportion of activated experts $\kappa=n_{\operatorname{act}}/n_{\operatorname{exp}}$. Our derivation also allows taking the expert size $n_{\operatorname{hid}}\triangleq\alpha_{\operatorname{ffn}}n_{\operatorname{embd}}\to\infty$ and the depth $L\to\infty$ so long as ${n_{\operatorname{embd}}}/({n_{\operatorname{exp}}n_{\operatorname{hid}}L)}$ is bounded in the limit. At a high level, our results reveal an asymptotic three-level mean-field structure made of residual stream neurons, expert outputs, and within-expert neurons:
@@
    The dynamics of the output $f$ is determined by a finite number of averages over residual stream neurons and over the state of experts within each hidden MoE layer.
@@
    The dynamics of each residual stream neuron depends on the rest of the network only through a finite number of averages over all other residual neurons and experts.
@@
    Sparse expert activation is determined by gating variables $q=\sigma(r)+b$ and set by a quantile threshold $q_{\star}(\kappa)$ which satisfies $\left[\mathbf{1}_{q\geq q_{\star}(\kappa)}\right]=\kappa$. The (hard) gating variables for each expert are thus $\sigma\triangleq\mathbf{1}_{q\geq q_{\star}(\kappa)}\sigma(r)$.
  4. 4.
    The dynamics of each expert (expert output, routing weights, and bias) depends on the rest of the network only through a finite number of averages over all other experts, all residual neurons, and all of its own neurons.
  5. 5.
    The dynamics of each neuron within an expert depends on the rest of the network only through a finite set of averages over all other neurons in the same expert, all other experts, and all residual neurons.
The averages in the preceding description obey deterministic evolution equations in the large $n_{\mathrm{embd}},n_{\mathrm{act}},L$ limit (see Appendix E for a full set of closed evolution equations for macroscopic variables). Several useful observations can be immediately gleaned from the theory:
@@
    $n_{\operatorname{hid}}$ does not need to diverge. If $n_{\operatorname{embd}},n_{\operatorname{exp}}\to\infty$, the output of the network obeys a deterministic evolution which depends on $n_{\operatorname{hid}}$, but not on $\nu\triangleq\frac{n_{\operatorname{exp}}}{n_{\operatorname{embd}}}$. While this limit is not studied empirically in our work, we point to (He, 2024) for an empirical investigation.
A similar three-level structure can be found for multi-head self-attention (Bordelon et al., 2024), where there exists a measure over attention heads (analogous to our measure over the experts) and an additional measure of key-query neurons within each head (analogous to our per-expert neurons). We also refer to Section 4 in (Bordelon et al., 2023) for a more thorough explanation of how mean-field limit of training dynamics supports reliable HP transfer.